OpenStudy (anonymous):

MEDAL What is the equation in point−slope form of the line passing through (2, 2) and (4, 8)

OpenStudy (jacob902):

irst of all, you need to know the point-slope form equation. point-slope form: y1 - y2 = m (x1 - x2) Then, find the slope of the line. slope equation: y1 - y2 / x1 - x2 8 - 3 / 2 - 3 5 / -1 -5 Next, plug in the coordinates in the right place. x2: 2 or 3 y2: 8 or 4 plug-in: y - 4 = -5 (x - 3) y - 8 = -5 (x - 2) And that's the equation you can use either coordinates if you'd like.

OpenStudy (anonymous):

(y+2)=3(x+2) should be the answer

OpenStudy (anonymous):

im i right

OpenStudy (ali2x2):

helped?

OpenStudy (ali2x2):

lemme do the math brb

OpenStudy (ali2x2):

yeah